Compute (a) return on equity (ROE), (b) market-to- book (M/B) ratio, and (c) enterprise value based on the following information: net profit margin = 4%, asset turnover = 1.5, debt = $60 million (assuming that the book value and market value of debt are the same), book value of assets = $100 million, market capitalization = $80 million, cash = $20 million.
Answer #1

(a) Total Assets = BV of Assets + Cash = 100+20 = 120 mn

Asset Turnover Ratio = Net Sales/Total Assets

i.e Net Sales/120 = 1.5

Therefore Net Sales = 120*1.5 = 180 mn

Net Profit = Net Profit Margin* Turnover = 4%*180 = 7. 2 mn

Shareholder's Equity = Total Assets - Total Debt = 120 - 60 = 60 mn

Therefore ROE = Net Profit/ Shareholder's Equity = 7.2/60 = 12%

(b) Market to book value = Market Cap/Net book value

Net book value = Total Assets - Total Liabilities

i.e. 120 - 60 = 60 mn

Therefore Market to book value = 80/60 = 1.33

(c) Enterprise Value = Market Cap +Total Debt - Cash

= 80+60-20

= 120 mn